OpenStudy (anonymous):

A bag contains 3 yellow tennis balls and 4 green tennis balls. Gordy will select one tennis ball, put it in his pocket, and then select a second tennis ball. Which expression gives the probability that he will select two green tennis balls? A. B. C. D.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

A. 4/7 . 3/6 B. 4/7 . 4/6 C. 4/7 . 3/7 D. 4/7 . 4/7

OpenStudy (anonymous):

c

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i will medal who answers first and says it is correct

OpenStudy (triciaal):

Event A and event B multiply the probabilities

OpenStudy (anonymous):

12

OpenStudy (triciaal):

4/7 * 3/6. Since not replaced

OpenStudy (anonymous):

4/7 . 3/7

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

or wrong

OpenStudy (triciaal):

Read my responses

OpenStudy (anonymous):

they r out of 7

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

why 3/6 not 3/7

OpenStudy (triciaal):

He kept the 1st so the 2nd is out of 6

OpenStudy (anonymous):

o k

OpenStudy (anonymous):

then it is 4/7 . 3/6

OpenStudy (anonymous):

is it

OpenStudy (anonymous):

Answer A 4/7 * 3/6 There are 3 + 4 = 7 tennis balls to start with and 4 of them are green. Therefore the probability of the first being green = 4/7. Once he puts this ball in his ocket there are only 6 balls left to select from and now only three are green. Remember the other green ball is in his pocket. Therefore the probability that the second is green also = 3/6. Finally as you are asked the probability that BOTH are green, multiply the individual probabilities together 4/7 * 3/6 = Answer A

OpenStudy (triciaal):

Yes. ( Use * for multiplication )

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

A volleyball team had 6 boys and 5 girls on it. The team needed 2 equipment managers. The coach selected 1 player at random to be an equipment manager and then selected 1 of the remaining players at random to be the other equipment manager. What is the probability that both equipment managers were girls? Express your answer in simplest form

OpenStudy (anonymous):

2/11

OpenStudy (anonymous):

2/11

OpenStudy (anonymous):

2/11 is correct

OpenStudy (anonymous):

stop searching it up u r in 6th

OpenStudy (triciaal):

When a question is posted anyone can answer. You do not decide who answers!! I do not doubt that you can answer or how good you are.

OpenStudy (anonymous):

you stay an hour just to type that wat kind of Q u r

OpenStudy (anonymous):

is it 2/!!

OpenStudy (anonymous):

weeeee know

OpenStudy (anonymous):

i unfaned you

OpenStudy (anonymous):

both ya

OpenStudy (anonymous):

no

OpenStudy (anonymous):

u turn

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@triciaal am i right or left

OpenStudy (anonymous):

2/11

OpenStudy (anonymous):

closing Q

OpenStudy (triciaal):

@Firedragon123 no at first 11 now only 10 to choose from. Key the remaining players to choose the 2nd manager

OpenStudy (anonymous):

5/11

OpenStudy (anonymous):

or 2/11

OpenStudy (anonymous):

5/11+4/11 what is left

OpenStudy (anonymous):

2/11

OpenStudy (anonymous):

closing the Q

OpenStudy (anonymous):

should i

OpenStudy (anonymous):

bye

OpenStudy (triciaal):

At first 11 total. 5 possible Next 10 total 1 girl already chosen only 4 girls left. So 4/10 Multiply the probabilities Reduce to the lowest term

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

5/10 - 4/10

OpenStudy (anonymous):

5/11

OpenStudy (triciaal):

5/11*4/10 = 2/11

OpenStudy (anonymous):

so 2/11 is correct

OpenStudy (triciaal):

Yes

OpenStudy (anonymous):

i said that in the beginning

OpenStudy (anonymous):

a moderator closed this Q

OpenStudy (anonymous):

bye passed

OpenStudy (anonymous):

u r making it more complicated

OpenStudy (triciaal):

So sorry you asked if you were correct. Will not answer any more of your questions.

OpenStudy (anonymous):

why

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

v